# Finite energy well-posedness for nonlinear Schrödinger equations with non-vanishing conditions at infinity

### v2

@@ Abstract.
Relevant physical phenomena are described by nonlinear Schrödinger equations with non-vanishing conditions at infinity. This paper investigates the respective 2D and 3D Cauchy problems. Local well-posedness in the energy space for energy-subcritical nonlinearities, merely satisfying Kato-type assumptions, is proven, providing the analogue of the well-established local $H^{1}$-theory for solutions vanishing at infinity. The critical nonlinearity will be simply a byproduct of our analysis and the existing literature. Under an assumption that prevents the onset of Benjamin-Feir instability, global well-posedness in the energy space is proven for a) non-negative Hamiltonians, b) sign-indefinite Hamiltonians under additional assumptions on the zeros of the nonlinearity, c) generic nonlinearities and small initial data. The cases b) and c) only concern 3D.
@@ 1. Introduction
and where the nonlinearity satisfies $f(\rho_{0})=0$ The Hamiltonian (coinciding with the total energy, in many relevant physical contexts) associated to (1.1)
@@
The finite energy assumption encodes (1.2) motivated by physical phenomena. Namely, we will deal with infinite energy solutions having finite relative energy with respect to the far-field state.
In the first part of the paper, our main purpose is to obtain a local well-posedness theory for (1.1) with energy-subcritical nonlinearities in a suitable energy space. We consider nonlinear potentials $f$ under Kato-type [39] regularity assumptions and we prove the continuity of the solution map with respect to the topology of the (curved) energy space and not only in affine spaces.
Global well-posedness is proven, provided that $f^{\prime}(1)>0$, see Assumption 1.5 below under different sets of assumptions. When $d=2,3$ for sign-definite total energies, when $d=3$ for sign-indefinite total energies by suitable additional assumptions on $f$ and some assumptions concerning the decay of data at infinity or in alternative for small initial data.
Regarding the $3D$-energy critical problem, we remark that global well-posedness is easily achieved relying on the existing literature [44, 18, 64] combined with our analysis for the sub-critical case.
Without loss of generality we assume (1.2) to hold for $\rho_{0}=1$ as the general case reduces to the former by a suitable scaling of $\psi$.
The most prominent example for (1.1), with far-field (1.2), is the Gross-Pitaevskii (GP) equation for which $f(\rho)=\rho-1$ arising in the description of Bose-Einstein condensates (BEC) [31, 59, 28, 60], as a model for superfluidity in Helium II close to the $\lambda$-point [27, 59] and for quantum vortices [59], see also the review paper [7] .
Beyond that, system (1.1)-(1.2) and general nonlinearities $f$ is relevant for the theory of (BEC), superconductivity and nonlinear optics. For instance, competing (focusing-defocusing) see e.g. (1.18), saturating or exponential nonlinearities emerge as models in nonlinear optics [5, 46, 51, 58]. Further physically relevant models are listed in Example 1.9 below.
The mathematical analysis of (1.1), with far-field behavior (1.2), differs significantly from the usual $H^{1}$-theory for NLS equations with trivial far-field, due to the non-integrability of the finite energy wave-functions, which may exhibit non-trivial oscillations at spatial infinity, in particular for $d=2$.
Opposite to the defocusing nonlinear Schrödinger equation with vanishing conditions at infinity for which scattering is known [26], system (1.1) with defocusing nonlinearity and equipped with (1.2) exhibits a very rich dynamics and admits a large variety of special solutions. Concerning GP, the existence of sub-sonic traveling waves is known for $d=2$ [10, 8] and $d=3$ [10, 9, 15], while non-existence in the super-sonic regime is proven in [29]. Traveling waves exist for arbitrarily small energy for $d=2$ [10], for $d=3$ non-existence of traveling waves, with small energy, is due to [8], see also [20] for $d\geq 3$.
@@ 1.1. Previous well-posedness results
To give a short overview of previous well-posedness results. Local existence of solutions to GP in Zhidkov spaces has been investigated in [66] for $d=1$, see also [68], and [21] for the multi-dimensional case. Global well-posedness of GP in an energy space is shown in [23] and for fractional Sobolev and low regularity for $d=1$ has recently been proven in [48, 49]. While the energy space for GP for $d=1$ coincides with the set of functions in the Zhidkov space, such that $|\psi|^{2}-1\in L^{2}({\mathbf{R}})$, this identification does not hold true in the multi-dimensional case, see [23] and Section 2 below. The GP is well-posed in $1+H^{1}({\mathbf{R}}^{d})$ for $d=2,3$ [10]. Global well-posedness in $1+H^{s}({\mathbf{R}}^{3})$ with $s\in(5/6,1)$ is proven in [57]. However, the space $1+H^{1}({\mathbf{R}}^{d})$ is strictly smaller than the natural energy space $\mathbb{E}({\mathbf{R}}^{d})$, see (1.9) below. In fact, there exist traveling waves for GP in the energy space that do not belong to $1+L^{2}({\mathbf{R}}^{d})$, see [29]. Global well-posedness in the energy space for the multi-dimensional GP has been introduced in the seminal paper [23]. One of the major novelties of [23] consists in the precise characterization of the energy space as complete metric space and the action of the free propagator on the energy space. A more general class of defocusing and energy-subcritical $C^{3}$-nonlinearities has been considered in [22] with subsequent improvement to $C^{2}$-nonlinearities [56]. In [22, 56], the respective authors crucially rely on a smooth decomposition of wave-functions in the energy space. Global well-posedness is proven in affine spaces determined by this decomposition which requires the aforementioned regularity assumptions and precise growth conditions for $f$. The result in the affine spaces then implies existence and uniqueness in the energy space. The cubic-quintic equation being energy-critical is studied in [44, 42, 43].
@@
Our first purpose is to prove local well-posedness assuming merely Kato-type regularity assumptions [39]. Further, the continuous dependence on the initial data is stated in the terms the topology of the energy space induced by its metric and not only in affine spaces with $H^{1}$-topology.
@@ 1.2. Local well-posedness results
These nonlinearities being included in Assumption 1.1 merely satisfy $f\in C^{0,\alpha}([0,\infty))$. It is straightforward to see that previous results require $\lambda=+1$ and $\alpha=1$ [23], $\lambda>0$, and $f\in C^{3}([0,\infty))$ [22] or $f\in C^{2}([0,\infty))$ [56]. The corresponding nonlinear potential energy density reads
@@
Note that (2) of Theorem 1.3 states that $\psi$ and $\psi_{0}$ share the same far-field behavior, i.e. belong to the same connected component of $\mathbb{E}({\mathbf{R}}^{d})$ for all $t\in[0,T^{\ast})$, see Remark 2.3 and 2.4. Moreover, it can be shown that the nonlinear flow $\psi-\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}$ belongs to the full range of Strichartz spaces, see Proposition 3.2 and 4.1 for $d=2,3$ respectively. The precise notion of continuous dependence on the initial data is given in Proposition 3.2 and 4.1. The topological structure of the metric space $(\mathbb{E}({\mathbf{R}}^{d}),d_{\mathbb{E}})$ differs for $d=2$ and $d=3$, see [23, 24]. For $d=3$, the energy space $\mathbb{E}({\mathbf{R}}^{3})$ has an affine structure; if $\psi\in\mathbb{E}({\mathbf{R}}^{3})$ then $\psi=c+v$ for some $c\in S^{1}$, $v\in\dot{H}^{1}({\mathbf{R}}^{3})$. For $d=2$, unbounded phase oscillations may occur at spatial infinity that rule out to characterize the connected components by a constant $c\in S^{1}$. The space $(\mathbb{E}({\mathbf{R}}^{2}),d_{\mathbb{E}})$ is not separable. Given its relevance for the well-posedness theory, this question is going to be addressed in detail in Section 2. In particular, one may introduce a weaker topology that restores separability and connectedness. Note that this affine structure of the energy space is available for higher dimensions $d\geq 4$ to which our approach adapts. As $\mathbb{E}({\mathbf{R}})\subset X^{1}({\mathbf{R}})$, the local well-posedness theory simplifies for $d=1$. Previous results [21, 22, 25] do not cover the full generality of Assumption 1.1. We expect our approach to extend to $d=1$.
Assumption 1.1 is not sufficient in order to prove that the solution map is Lipschitz continuous. This is analogue to the case of NLS equations (1.1) with vanishing far-field behavior. Indeed, for instance for power-law type nonlinearities (1.4) Lipschitz continuity of the solution map can only be expected if $\alpha\geq\frac{1}{2}$ for both vanishing and non-vanishing far-field, see [14, Remark 4.4.5] and Section 5 respectively.
@@
(1.14) $f\in C^{1}([0,\infty))\cap C^{2}((0,\infty)),\qquad\left|\sqrt{\rho}f^{\prime}(\rho)\right|,\left|\rho^{\frac{3}{2}}f^{\prime\prime}(\rho)\right|\leq C(1+\rho^{\max\{0,\alpha-\frac{1}{2}\}}),$
@@ 1.3. Global well-posedness results
This assumption yields that $F$ achieves a local minimum for the constant solution $|\psi|^{2}=1$. In nonlinear optics, this requirement is made in physical literature in order to prevent the onset of modulational instability, also referred to as Benjamin-Feir instability [6], of the constant equilibrium solution, i.e. the continuous wave background [45, 58]. On the level of the corresponding quantum hydrodynamic system (QHD), Assumption 1.5 ensures the hyperbolicity of QHD linearized around the equilibrium $(\rho=\rho_{0},v=0)$ neglecting the quantum correction.
@@
The assumption on the roots of $F$ allows for physically relevant nonlinearities to be studied. It appears from the physics literature [5, 46, 58] that in relevant applications the largest root of $F$ corresponds to the far-field behavior $\rho_{0}=1$ and constitutes a local minimum of $F$ which is linked to the modulational stability of the continuous background wave [45, 58]. To obtain global existence, we rely on the aforementioned affine structure of the energy space $\mathbb{E}({\mathbf{R}}^{3})$ and require that $\operatorname{Re}(v_{0})\in L^{2}({\mathbf{R}}^{3})$ while $v_{0}\in\mathcal{F}_{c}({\mathbf{R}}^{3})$ only yields $|v_{0}|^{2}+2\operatorname{Re}(\overline{c}v_{0})\in L^{2}({\mathbf{R}}^{2})$. An exponential bound on $\operatorname{Re}(\overline{c}v)(t)\in L^{2}({\mathbf{R}}^{3})$ is derived which compensates for the lack of the conserved mass due to the non-trivial farfield. The result of Theorem 1.7 remains valid if the assumption $\operatorname{Re}(\overline{c}v)\in L^{2}$ is replaced by a smallness assumption on $\mathcal{H}(\psi_{0})$ and $\|\nabla\operatorname{Re}(\overline{c}v_{0})\|_{L^{2}}^{2}$ depending only on the second largest positive root of $F$, see Remark 4.12.
@@
As it will become clear from the proofs, our method adapts to prove well-posedness for energy-sub-critical nonlinearities for $d\geq 4$. For the energy-critical critical quintic equation, one may proceed as described in Section 1.4.
@@ 4.1. Global well-posedness
Under Assumption 4.7 and instead of $\operatorname{Re}(v_{0})\in L^{2}({\mathbf{R}}^{3})$, one may alternatively assume that the initial data satisfies $\mathcal{H}(1+v_{0})$ and $\|\nabla\operatorname{Re}(v_{0})\|_{L^{2}}^{2}$ sufficiently small adapting [42, Lemma 3.2] stated for cubic-quintic nonlinearities (1.18). Moreover, as pointed out in [43, Remark p. 2683] the same argument yields small data global well-posedness for the cubic-quintic nonlinearity where the quintic part is focusing and the cubic part defocusing, hence for $F$ being unbounded from below. Inspired, by this observation and the classical small data global well-posedness in $H^{1}$ for NLS eq, see e.g. [62, Chapter 3.4], we prove that (4.3) is globally well-posed in the energy space for small data provided that Assumption (1.5) holds.
@@
If Assumptions 1.5 are satisfied, then there exists $\varepsilon>0$ only depending on $\delta>0$ as in (2.17) such that if $\mathcal{H}(1+v_{0})\leq\frac{1}{4}\varepsilon$ and $\|\nabla v_{0}\|_{L^{2}}^{2}\leq\varepsilon$, then the unique solution provided by Proposition 4.1 is global.
@@ Acknowledgments
L.E.H. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 317210226 – SFB 1283.

### v3

@@ Abstract.
Relevant physical phenomena are described by nonlinear Schrödinger equations with non-vanishing conditions at infinity. This paper investigates the respective 2D and 3D Cauchy problems. Local well-posedness in the energy space for energy-subcritical nonlinearities, merely satisfying Kato-type assumptions, is proven, providing the analogue of the well-established local $H^{1}$-theory for solutions vanishing at infinity. The critical nonlinearity will be simply a byproduct of our analysis and the existing literature. Under an assumption that prevents the onset of a Benjamin-Feir type instability, global well-posedness in the energy space is proven for a) non-negative Hamiltonians, b) sign-indefinite Hamiltonians under additional assumptions on the zeros of the nonlinearity, c) generic nonlinearities and small initial data. The cases b) and c) only concern 3D.
@@ 1. Introduction
and where the nonlinearity satisfies $f(\rho_{0})=0$. Without loss of generality, we assume $\rho_{0}=1$ as the general case is obtained by a suitable scaling. The Hamiltonian (coinciding with the total energy, in many relevant physical contexts) associated to (1.1) is given by
@@
The finite energy assumption encodes (1.2). Namely, we deal with infinite energy solutions having finite relative energy with respect to the far-field state.
The system (1.1)-(1.2) appears in relevant physical applications. Most prominently, the Gross-Pitaevskii (GP) equation, i.e. $f(\rho)=\rho-1$, is studied as model for Bose-Einstein condensates (BEC) [31, 59, 28, 60], superfluidity in Helium II close to the $\lambda$-point [27, 59] and for quantum vortices [59], see also [7]. Competing (focusing-defocusing) see e.g. (1.18), saturating or exponential nonlinearities for (1.1)-(1.2) emerge as models in nonlinear optics [5, 46, 51, 58]. Further physically relevant models are listed in Example 1.9 below.
In the first part of the paper, we establish local well-posedness in the energy space for (1.1)-(1.2) with energy-subcritical nonlinear potentials $f$ under Kato-type [39] regularity assumptions. The continuity of the solution map is proven with respect to the topology of the (curved) energy space and not only in affine spaces.
Second, global well-posedness is proven, provided that $f^{\prime}(1)>0$, see Assumption 1.5 below. Specifically, global well-posedness is shown for sign-definite total energies and $d=2,3$, and for sign-indefinite total energies and $d=3$ under suitable additional assumptions on $f$ and the decay of the initial data at infinity or in alternative for small initial data.
Regarding the $3D$-energy critical problem, we remark that global well-posedness is easily achieved relying on the existing literature [44, 18, 64] combined with our analysis for the sub-critical case, see Section 1.4.
The mathematical analysis of (1.1), with far-field behavior (1.2), differs significantly from the usual $H^{1}$-theory for NLS equations with trivial far-field. Finite energy wave-functions are not integrable and may exhibit non-trivial oscillations at spatial infinity, in particular for $d=2$.
System (1.1)-(1.2) with defocusing nonlinearity exhibits a very rich dynamics and admits a large variety of special solutions, contrary to the case of vanishing far-field [26]. Concerning the GP equation, the existence of sub-sonic traveling waves is known for $d=2$ [10, 8] and $d=3$ [10, 9, 15], while non-existence in the super-sonic regime is proven in [29]. Traveling waves exist for arbitrarily small energy for $d=2$ [10]. On the contrary, for $d=3$ non-existence of traveling waves with small energy is proven in [8, 20].
@@ 1.1. Previous well-posedness results
Local existence of solutions to GP in Zhidkov spaces has been investigated in [66, 68] for $d=1$ and [21] for the multi-dimensional case. For $d=1$, global well-posedness of GP in the energy space is shown in [66] and has recently been proven for fractional Sobolev and low regularity in [48, 49]. While the energy space for GP for $d=1$ coincides with the set of functions in the Zhidkov space, such that $|\psi|^{2}-1\in L^{2}({\mathbf{R}})$, this identification does not hold true in the multi-dimensional case, see [23] and Section 2 below. The GP is well-posed in $1+H^{1}({\mathbf{R}}^{d})$ for $d=2,3$ [10]. Global well-posedness in $1+H^{s}({\mathbf{R}}^{3})$ with $s\in(5/6,1)$ is proven in [57]. However, the space $1+H^{1}({\mathbf{R}}^{d})$ is strictly smaller than the natural energy space $\mathbb{E}({\mathbf{R}}^{d})$, see (1.9) below. In fact, there exist traveling waves for GP in the energy space that do not belong to $1+L^{2}({\mathbf{R}}^{d})$, see [29]. Global well-posedness in the energy space for the multi-dimensional GP has been introduced in the seminal paper [23]. One of the major novelties of [23] consists in the precise characterization of the energy space as complete metric space and the action of the free propagator on the energy space. A more general class of defocusing and energy-subcritical $C^{3}$-nonlinearities has been considered in [22] with subsequent improvement to $C^{2}$-nonlinearities [56]. In [22, 56], the respective authors crucially rely on a smooth decomposition of wave-functions in the energy space. Global well-posedness is proven in affine spaces determined by this decomposition which requires the aforementioned regularity assumptions and precise growth conditions for $f$. The result in the affine spaces then implies existence and uniqueness in the energy space. The cubic-quintic equation being energy-critical is studied in [44, 42, 43].
@@
1.2. Local well-posedness results
Our first purpose is to prove local well-posedness assuming merely Kato-type regularity assumptions [39] and with the continuous dependence on the initial data is stated with respect to the topology of the energy space.
@@ 1.2. Local well-posedness results
These nonlinearities being included in Assumption 1.1 merely satisfy $f\in C^{0,\alpha}([0,\infty))$. Previous results require $\lambda=+1$ and $\alpha=1$ [23], $\lambda>0$, and $f\in C^{3}([0,\infty))$ [22] or $f\in C^{2}([0,\infty))$ [56]. The corresponding nonlinear potential energy density reads
@@
Note that (2) of Theorem 1.3 states that $\psi$ and $\psi_{0}$ share the same far-field behavior, i.e. belong to the same connected component of $\mathbb{E}({\mathbf{R}}^{d})$ for all $t\in[0,T^{\ast})$, see Remark 2.3 and 2.4. Moreover, it can be shown that the nonlinear flow $\psi-\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}$ belongs to the full range of Strichartz spaces, see Proposition 3.2 and 4.1 for $d=2,3$ respectively. The precise notion of continuous dependence on the initial data is given in Proposition 3.2 and 4.1. The topological structure of the metric space $(\mathbb{E}({\mathbf{R}}^{d}),d_{\mathbb{E}})$ differs for $d=2$ and $d=3$, see [23, 24]. For $d=3$, the energy space $\mathbb{E}({\mathbf{R}}^{3})$ has an affine structure; if $\psi\in\mathbb{E}({\mathbf{R}}^{3})$ then $\psi=c+v$ for some $c\in S^{1}$, $v\in\dot{H}^{1}({\mathbf{R}}^{3})$. For $d=2$, unbounded phase oscillations may occur at spatial infinity that rule out to characterize the connected components by a constant $c\in S^{1}$. The space $(\mathbb{E}({\mathbf{R}}^{2}),d_{\mathbb{E}})$ is not separable. Given its relevance for the well-posedness theory, this question is going to be addressed in detail in Section 2. In particular, one may introduce a weaker topology that restores separability and connectedness. Note that this affine structure of the energy space is available for higher dimensions $d\geq 4$ to which our approach adapts. As $\mathbb{E}({\mathbf{R}})\subset X^{1}({\mathbf{R}})$, the local well-posedness theory simplifies for $d=1$. We expect our approach to extend to $d=1$. Previous results [21, 22, 25] do not cover the full generality of Assumption 1.1.
Assumption 1.1 is not sufficient in order to prove that the solution map is Lipschitz continuous. This is analogue to the $H^{1}$-theory for (1.1) with vanishing far-field behavior. Indeed, for instance for power-law type nonlinearities (1.4) Lipschitz continuity of the solution map can only be expected if $\alpha\geq\frac{1}{2}$ for both vanishing and non-vanishing far-field, see [14, Remark 4.4.5] and Section 5 respectively.
@@
(1.14) $f\in C^{1}([0,\infty))\cap C^{2}((0,\infty)),\,\,\left|\sqrt{\rho}f^{\prime}(\rho)\right|,\left|\rho^{\frac{3}{2}}f^{\prime\prime}(\rho)\right|\leq C(1+\rho^{\max\{0,\alpha-\frac{1}{2}\}}),$
@@ 1.3. Global well-posedness results
This assumption yields that $F$ achieves a local minimum for the constant solution $|\psi|^{2}=1$. In nonlinear optics, this requirement is made in physical literature in order to prevent the onset of modulational instability, also referred to as Benjamin-Feir instability [6], of the constant equilibrium solution, i.e. the continuous wave background [45, 58].
@@
The assumption on the roots of $F$ allows for physically relevant nonlinearities to be studied. It appears from the physics literature [5, 46, 58] that in relevant applications the largest root of $F$ corresponds to the far-field behavior $\rho_{0}=1$ and constitutes a local minimum of $F$ which is linked to preventing modulational instability of the continuous background wave [45, 58]. To obtain global existence, we rely on the aforementioned affine structure of the energy space $\mathbb{E}({\mathbf{R}}^{3})$ and require that $\operatorname{Re}(\overline{c}v_{0})\in L^{2}({\mathbf{R}}^{3})$ while $v_{0}\in\mathcal{F}_{c}({\mathbf{R}}^{3})$ only yields $|v_{0}|^{2}+2\operatorname{Re}(\overline{c}v_{0})\in L^{2}({\mathbf{R}}^{2})$. An exponential bound on $\operatorname{Re}(\overline{c}v)(t)\in L^{2}({\mathbf{R}}^{3})$ is derived which compensates for the lack of the conserved mass due to the non-trivial farfield. The result of Theorem 1.7 remains valid if the assumption $\operatorname{Re}(\overline{c}v_{0})\in L^{2}$ is replaced by a smallness assumption on $\mathcal{H}(\psi_{0})$ and $\|\nabla\operatorname{Re}(\overline{c}v_{0})\|_{L^{2}}^{2}$ depending only on the second largest positive root of $F$, see Remark 4.12.
@@
For the 3D energy-critical critical quintic equation, one may proceed as described in Section 1.4.
@@ 4.1. Global well-posedness
Under Assumption 4.7 and instead of $\operatorname{Re}(v_{0})\in L^{2}({\mathbf{R}}^{3})$, one may alternatively assume that the initial data satisfies $\mathcal{H}(1+v_{0})$ and $\|\nabla\operatorname{Re}(v_{0})\|_{L^{2}}^{2}$ sufficiently small adapting [42, Lemma 3.2] stated for cubic-quintic nonlinearities (1.18). Moreover, as pointed out in [43, Remark p. 2683] the same argument yields small data global well-posedness for the cubic-quintic nonlinearity where the quintic part is focusing and the cubic part defocusing, hence for $F$ being unbounded from below. Inspired, by this observation and the classical small data global well-posedness in $H^{1}$ for NLS eq, see e.g. [62, Chapter 3.4], we prove that (4.3) is globally well-posed in the energy space for small data provided that Assumption 1.5 holds.
@@
If Assumption 1.5 is satisfied, then there exists $\varepsilon>0$ only depending on $\delta>0$ as in (2.17) such that if $\mathcal{H}(1+v_{0})\leq\frac{1}{4}\varepsilon$ and $\|\nabla v_{0}\|_{L^{2}}^{2}\leq\varepsilon$, then the unique solution provided by Proposition 4.1 is global.
@@ Acknowledgments
P.A. and P.M. acknowledge partial support from INdAM-GNAMPA. P.A. is partially supported by PRIN project 20204NT8W4 ”Nonlinear evolution PDEs, fluid dynamics and transport equations: theoretical foundations and applications”. L.E.H. is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 317210226 – SFB 1283.